# LLMs in the SOC: An Empirical Study of Human-AI Collaboration in Security Operations Centres

### v1

@@ 4.2.2 SOC Analyst Query Subjects
Telemetry tells us what was observed; analyst queries reveal which aspects they seek to interpret using LLMs. While security teams collect rich telemetry via tools like SIEMs and EDRs, LLMs are increasingly used to interpret parts of that raw input, to potentially transform the raw input into actionable insights or contextual understanding. Analysts appear to have relied on LLMs to interpret telemetry-derived artifacts (e.g., commands, log fragments), to likely improve situational awareness and support timely decision-making.
The MITRE ATT&CK data sources [^2] https://attack.mitre.org/datasources/ describe telemetry collected by sensors (e.g., process creation, file access, authentication logs). Most analyst queries focused on interpreting artifacts that these sensors collect, such as commands, scripts, OS processes or file paths. While subject themes do not directly map to MITRE ATT&CK data sources, they reflect reasoning tasks shaped by this telemetry. For instance, when an analyst asked the LLM to explain a PowerShell command launching a scheduled task, that artifact maps to ‘Command Execution’.
@@ 4.2.4 Alignment with Professional Cybersecurity Frameworks
Building on the task and query subject, we explored whether analyst queries aligned with established cybersecurity roles and competencies. We reviewed all queries against the NICE Framework for Cybersecurity [^3] https://www.nist.gov/itl/applied-cybersecurity/nice/nice-framework-resource-center by matching each task and subject to the NICE categories.
This comparison revealed that 93% of analyst queries could be reasonably associated with at least one NICE Task, Knowledge, or Skill element; the other 7% mostly included task themes like LLM Chat or Other. For instance, command interpretation queries align with K0805 (Knowledge of command-line tools and techniques), while email crafting aligns with S0610 (effective communication skills).
@@ 4.3 Conversation Level Analysis
Our thematic coding centred on analyst queries, aligning with our objective of understanding how SOC analysts engage with LLMs in practice. Although we reviewed LLM responses to contextualise conversation-level analyst intent, we did not formally code them. This decision was based on the following considerations: (1) LLM outputs can vary in content or structure even when prompted identically, depending on deployment settings, making systematic analysis less reliable; (2) our primary research objective was to analyse analyst behaviour rather than audit or benchmark the LLM itself; and (3) defining response utility is highly analyst- and context-dependent and often lacks objective ground truth. This analytic focus allowed us to prioritise the human perspective, capturing analyst reasoning patterns and interaction strategies, without over-attributing meaning to the LLM’s behaviour, which may vary depending on deployment conditions/settings or model version.
@@ 4.3.1 Conversation Lengths
Table I summarises conversation length and the associated mean and median durations (elapsed time between first and last analyst message; LLM response time is not measured). Most conversations were brief: 57% involved just two analyst steps, and 75% contained only 2–3 queries. Only a small fraction (just over 4%) exceeded 10 steps. Durations increase with length, but longer conversations show much higher variance, with mean durations often far exceeding the median, suggesting that many multi-turn sessions are prolonged by idle periods or interruptions rather than sustained engagement. The following sections unpack how tasks and reasoning patterns unfold across these conversation lengths.
TABLE I: Summary of Conversation Lengths and Durations
@@ 4.3.2 Most Common Two-Step Conversation Sequences
Analysis of the most frequent two-step task sequences reveals distinct usage patterns in analyst-LLM interactions. The most dominant 2-step sequence involves back-to-back instances of Command Understanding/Analysis, which occur in 16.2% (of all conversations). This suggests that analysts often explore multiple related commands in succession or seek clarification on different aspects of the same command context. In these conversations, analysts commonly asked variations of “What does [x] do?”, “Explain this [x]” and “Can you explain what [x] does?” about a command in both steps. Another recurring sequence is Command Understanding/Analysis followed by Summarising Command (3% instances; in the top 5 of all sequences), reflecting a typical workflow where analysts first understand a command and then summarise it, likely for reporting purposes.
Several editing-related sequences appear frequently. These include repeated instances of Editing/Rewriting Text (11% instances), and combinations indicating that analysts often revise content iteratively with a focus on tone or clarity. The third most frequent loop (4.7%) is repeated Code, Script & Regex Analysis, reflecting multiple passes of script writing or debugging. Several other two-step patterns (1-3% of conversations) include self-loops in Attack & Threat Intel, Tool/Software Understanding, Query/Search & Data Processing, Network Analysis, and Email Analysis.
@@ 4.3.3 Three-Step Conversation Sequences
Among conversations consisting of three analyst steps, we observe multiple instances where analysts remain focused on a single task theme throughout. For example, we observed [Editing/Rewriting Text, Editing/Rewriting Text, Editing/Rewriting Text] sequences, where analysts revised text across multiple iterations (5%). These included sequential refinements such as “Make this better [x]”, followed by “Make it shorter [x]” and then “Make it sound better [x]”. Such sequences illustrate how analysts leverage the LLM as a drafting and refinement tool, shaping content, such as alerts, investigation notes, or emails, over several small adjustments.
There were triplets of command understanding-related queries (4%), typically involving repeated or slightly varied versions of prompts such as “What does [x] do?” or “Explain [x]”. Similar to earlier discussions, this reveals a pattern of iterative probing, where analysts utilise LLMs to interpret one or more commands or their components.
Interestingly, a small number of sequences involved transitions between different task themes (e.g., from command analysis to text editing, or alternating between command and tool understanding), highlighting cases where the analyst first interprets a command and then translates that interpretation/understanding into an incident narrative.
Overall, these three-step interactions reveal how analysts engage the LLM to incrementally make sense of complex information, refine their articulation, and support analytical workflows. The sequences reflect real-world cognitive processes, particularly when dealing with uncertainty or unfamiliar commands or scenarios.
@@ 4.3.4 Longer Conversations
While shorter interactions (lengths 2-3) were most common in our dataset, several conversations extended beyond three exchanges. Although individually less frequent, these longer interactions offer insight into how analysts build upon prior queries to progressively deepen understanding, integrate multiple concepts, or iteratively refine their outputs.
These longer conversations spanned a wide range of tasks. Some involved exploratory learning or upskilling, such as understanding machine learning fundamentals, understanding authentication mechanisms, or the Remote Desktop Protocol. Others reflected hands-on problem-solving, such as writing or debugging scripts, analysing remote shell behaviour, analysing registry modifications, or investigating potentially malicious behaviour (e.g., detecting potential ransomware). A distinct subset of longer conversations centred on professional development, likely related to CVs or performance reviews. These cases highlight the LLM’s use beyond core SOC functions, supporting analysts in broader professional contexts.
Many extended interactions show a progression from low-level technical understanding to higher-order reasoning. For example, conversations with four or more consecutive Command Understanding/Analysis steps suggest thorough exploration and careful reasoning of related commands in high-stakes environments where misinterpreting command behaviour could lead to erroneous threat assessments. Other sequences chain distinct but related subtasks, combinations such as File Understanding $\rightarrow$ Command Understanding/Analysis $\rightarrow$ Threat Analysis $\rightarrow$ Summarise LLM Output indicate layered workflows in which an analyst first clarifies data sources or command behaviour, then assesses threat implications, and finally requests help formulating an incident report or alert. We also observed iterative development conversations with repeated cycles of Code Writing and Understanding, Understanding Errors, and Query Writing, suggesting a debugging or prototyping mindset in which the LLM supports implementation, error checking, refinement, and formatting. Finally, several multi-step interactions centred on communication tasks, Editing/Rewriting Text, Improving Phrasing/Clarity, and Generating Empathetic Emails, combine structural edits with tone adjustment and simplification, highlighting the LLM’s utility in producing polished, context-appropriate communication for internal and external stakeholders.
Across all conversation lengths, a picture emerges of how SOC analysts engage with LLMs: as iterative, context-sensitive tools for reasoning, drafting, and refinement. Analysts often employ multiple steps to make sense of complex information and/or refine outputs. They use LLMs in deliberate, task-oriented ways, typically through short, focused exchanges that involve probing for interpretation, refining written outputs, or shifting between related tasks (from sense-making to writing).
@@ 4.4 Analyst Level Analysis
Our analysis reveals substantial variation in how individual analysts engaged with the LLM, both in terms of query volume and the thematic breadth of their interactions. As noted previously, analysts were grouped into three distinct clusters: low-usage analysts (Cluster 0, $n=34$), moderate users (Cluster 1, $n=10$), and a single high-volume outlier (Cluster 2, $n=1$; most active analyst).
We examined daily LLM query activity across all analysts (Figure 6) and observed distinct engagement patterns. First, one analyst (A68f) used the system consistently and at high volume across multiple months. Second, a group of moderate users (e.g., Ac9a, A5b3) engaged in concentrated bursts of activity, likely linked to specific investigations. In contrast, most analysts issued only a handful of queries on scattered days, suggesting sporadic or task-specific usage. These behavioural differences support cluster-based segmentation and reveal diverse styles of LLM integration within the analyst population.
@@
A30d and A5b3, while heavily focused on Code, Script, and Regex Analysis, did explore other tasks moderately. Adf2 showed a more balanced profile, spanning Text Generation, Command Understanding, and Tool/System Understanding, reflecting the use of the LLM for both exploratory and functional support tasks. Analysts like A4df combined command and code-related tasks with moderate Query/Data Processing and File Analysis, suggesting use cases grounded in technical inspection.
Among lower-volume users, such as A180, Ac48, and A95e, task engagement was more narrowly focused, typically confined to one or two functional categories such as Text Editing or Command Understanding. Notably, the recurring appearance of Command Understanding and Text Generation across many analysts underscores their central role in LLM support within SOC workflows.
@@
We observed that LLMs are not used uniformly across analysts. Instead, individuals adapt the tool to their task demands, specific needs, and investigative styles, ranging from quick clarification and script debugging to collaborative drafting and iterative refinement. The LLM successfully supported this diversity without requiring extensive prompt engineering. This highlights that a well-configured, domain-aligned LLM can support a broad range of task types and prompt styles, enabling quick and effective interaction, even without deep prompting expertise.
@@ 4.4.1 Discussion on Potential Disengagement Patterns
A2f7 repeatedly queried the model version (perhaps unmet transparency need), A5a6 received two consecutive erroneous answers, with a misclassification ending the session, and A6e9 disengaged immediately after the LLM refused a request. Another analyst (Ae46) ceased engagement after several command analysis queries, although no clear failure pattern was identified. Collectively, this group suggests that early disengagement occurred for multiple reasons: in some cases without clear frustration, and in others following specific model failures.
Group 2: Persistent but Mixed-Experience Analysts (n = 6) These analysts remained active for four to eight months. Two analysts (Ae32, Aae4) experienced apparent errors yet persisted, indicating that perceived utility can offset isolated failures. Other patterns suggest diverse usage motives: A9c2 moved past an early debugging issue; A214 relied on the LLM for repeated Sumo Logic refinements; A8bb may have engaged only when necessary; and A525 mixed non-SOC queries with SOC ones. This group suggests that analysts tolerate occasional mistakes, perhaps if using the LLM for specific tasks is beneficial.
While these patterns offer insights, we did not have direct access to analysts (due to privacy reasons) to confirm these. Future work should involve observing SOC analysts in their environment and conducting follow-up interviews to validate reasons for disengagement. This will be essential as LLM-based capabilities continue to evolve in SOCs.
@@ 5 Discussion
This study set out to investigate a core question (restated briefly): How do SOC analysts utilise LLMs in their daily workflows, including the specific tasks they apply them to, how these uses align with established cybersecurity frameworks, and the conversational patterns that characterise their interactions? Below, we address this question through a multi-level analysis and conclude with a discussion of the study’s limitations and the broader implications for designing LLM-based assistance in SOCs.
@@ 5.1 Query Level: Analyst Needs and Task Focus
The progressive increase in LLM use across different tasks (Figure 6) and sustained engagement by a growing number of analysts (Figure 3) indicate that analysts incrementally integrated the LLM into their daily workflows, applying it to an increasing number of tasks (see Figure 9 in Appendix for task adoption over time). These usage patterns suggest that for some analysts the LLM delivered sufficient value to merit repeated use during investigative sessions.
@@
Query patterns reflected clear cognitive intent. Analysts’ queries were not just syntactic variations but reflected clear cognitive intent. Functional queries such as “what does [x] do” consistently mapped to interpretive tasks like command understanding, while requests like “reword this” signalled refinement goals tied to documentation. This high alignment between linguistic pattern and task type suggests that query structure can act as a proxy for the analyst’s immediate reasoning objective.
@@ 5.3 Analyst-Level: Usage Patterns and Integration Approaches
LLM use varies. Analysts differed not only in the frequency of LLM engagement but also in the types of tasks they delegated and how they framed their requests. Some analysts primarily used the LLM for command interpretation and technical debugging, while others focused more narrowly on editing or summarisation tasks. These distinctions suggest an opportunity for investigating the effectiveness of role-aware, adaptive integration strategies that personalise tooling based on emerging usage patterns. Our findings extend early claims about LLMs as cognitive aides in SOC workflows [7, 8], and validate emerging theoretical framings that emphasise diverse user strategies and interaction modes [13, 35]. While prior work has focused on how LLMs can support structured information extraction or task acceleration, we demonstrate how usage patterns naturally emerge around analysts’ task orientations without the need for prompt engineering or adaptation.
LLMs function primarily as evidence interpreters, not decision-makers. Only 4% of queries sought explicit recommendations (e.g., “is this malicious?”). Instead, most analysts requested explanation, interpretation, or contextualisation, highlighting a strong preference for maintaining decision authority. This supports the design of LLMs as evaluative cognitive aides that surface relevant evidence rather than issuing prescriptive judgments. Our findings resonate with and provide support for broader concerns about over-reliance on automated suggestions [36, 37], reinforcing the need for machine-in-the-loop systems that enhance but do not override expert judgment [39, 40, 42]. We find that most analysts favour evidence interpretation over explicit recommendations. This opens doors for further research into why different analysts prefer different modes and whether there is a notable impact on their performance.
Other factors that modulate routine engagement: isolated errors and trust.
@@
Research Opportunity: Studies could investigate whether the preference for evidence-oriented outputs (4% in our study) generalise to different settings and SOC, and also compare evidence-oriented versus recommendation-oriented outputs on accuracy, response time, and calibrated trust [36, 37] to determine which support style minimises misuse yet maintains efficiency during high-pressure triage.

### v2

@@ 4.3.2 SOC Analyst Query Subjects
Telemetry tells us what was observed; analyst queries reveal which aspects they seek to interpret using LLMs. While security teams collect rich telemetry via tools like SIEMs and EDRs, LLMs are increasingly used to interpret parts of that raw input, transforming it into potentially actionable insights or contextual understanding. Analysts appear to have relied on LLMs to interpret telemetry-derived artifacts (e.g., commands, log fragments), to likely improve situational awareness and support timely decision-making.
The MITRE ATT&CK data sources [^5] https://attack.mitre.org/datasources/ describe telemetry collected by sensors (e.g., process creation, file access, authentication logs). Most analyst queries focused on interpreting artifacts that these sensors collect, such as commands, scripts, OS processes or file paths. While subject themes do not directly align with MITRE ATT&CK data sources, they reflect reasoning shaped by telemetry. For example, a PowerShell command launching a scheduled task maps to Command Execution.
@@ 4.3.4 Alignment with Professional Cybersecurity Frameworks
Building on the task and query subject, we explored whether analyst queries aligned with established cybersecurity roles and competencies. We reviewed all queries against the NICE Framework for Cybersecurity [^6] https://www.nist.gov/itl/applied-cybersecurity/nice/nice-framework-resource-center by matching each task and subject to the NICE categories. This comparison revealed that 93% of analyst queries could be reasonably associated with at least one NICE Task, Knowledge, or Skill element; the other 7% mostly included task themes like LLM Chat or Other. For instance, command interpretation queries align with K0805 (Knowledge of command-line tools and techniques), while email crafting aligns with S0610 (effective communication skills).
@@ 4.4 Conversation Level Analysis
One-off interactions comprised 41% of all interactions; the remaining 59% formed 532 multi-step conversations. The task themes ordering was stable across interaction types: Command Understanding/Analysis was most common, followed by Text Processing, then Code/Script tasks.
@@
Our thematic coding centred on analyst queries, aligning with our objective of understanding how SOC analysts engage with LLMs in practice. Although we reviewed LLM responses to contextualise conversation-level analyst intent, we did not formally code them. This decision was based on the following considerations: (1) LLM outputs can vary in content or structure even when prompted identically, depending on deployment settings, making systematic analysis less reliable; (2) our primary research objective was to analyse analyst behaviour rather than audit or benchmark the LLM itself; 3) defining response utility is highly analyst- and context-dependent and often lacks objective ground truth; and (4) LLM responses were de-identified, making it difficult to judge the full task context. This analytic focus allowed us to prioritise the human perspective, capturing analyst reasoning patterns and interaction strategies, without over-attributing meaning to the LLM’s behaviour, which may vary depending on deployment or model settings.
@@ 4.4.1 Conversation Lengths
Table II summarises conversation length and the associated mean and median durations (elapsed time between first and last analyst message; LLM response time is not measured). Most conversations were brief: 57% involved just two analyst steps, and 75% contained only 2–3 queries. Only a small fraction (just over 4%) exceeded 10 steps. Durations increase with length, but longer conversations show much higher variance, with mean durations often far exceeding the median, suggesting that many multi-turn sessions are prolonged by idle periods or interruptions rather than sustained engagement. The following sections unpack how tasks and reasoning patterns unfold across these conversation lengths.
TABLE II: Summary of Conversation Lengths and Durations
@@ 4.4.2 Most Common Two-Step Conversation Sequences
Analysis of frequent two-step sequences shows distinct patterns, with the most common being consecutive Command Understanding/Analysis (16.2% of conversations). This suggests that analysts often explore multiple related commands in succession or seek clarification on different aspects of the same command. Analysts often asked variations of “What does [x] do?” in both steps. Another common sequence is Command Understanding/Analysis followed by Summarising Command (3%), reflecting a workflow of interpreting and then summarising commands for reporting.
Several editing-related loops include repeated Editing/Rewriting Text (11%) and iterative revisions for tone or clarity. The third most common loop (4.7%) is repeated Code, Script & Regex Analysis. Other two-step patterns (1–3%) involve Attack & Threat Intel, Tool/Software Understanding, Query/Search Related, and Email Analysis.
@@ 4.4.3 Three-Step Conversation Sequences
In three-step conversations, analysts often stayed on one theme, e.g., [Editing/Rewriting Text, Editing/Rewriting Text, Editing/Rewriting Text] (5%), making iterative refinements like “Make this better”, then “Make it shorter”, and “Make it sound better”, showing that the LLM was used as a drafting and refinement tool for alerts, notes, or emails. Triplets of command understanding queries (4%) often repeated prompts like “What does [x] do?” or “Explain [x]”, showing iterative probing to interpret commands or components.
A small number of sequences involved transitions between different task themes (e.g., from command analysis to text editing, or alternating between command and tool understanding), highlighting cases where the analyst first interprets a command and then translates that understanding into an incident narrative. These three-step interactions reveal how analysts engage the LLM to incrementally make sense of complex information, refine their articulation, and support analytical workflows. The sequences reflect real-world cognitive processes, particularly when dealing with uncertainty or unfamiliar commands or scenarios.
@@ 4.4.4 Longer Conversations
Most interactions were short (2–3 turns), but longer ones, though rarer, show how analysts build on prior queries to deepen insight and refine results.
These longer conversations spanned a wide range of tasks. Some involved exploratory learning or upskilling, such as understanding machine learning fundamentals or understanding authentication mechanisms. Others reflected hands-on problem-solving, such as writing or debugging scripts, analysing remote shell behaviour, analysing registry modifications, or investigating potentially malicious behaviour (e.g., detecting potential ransomware). A distinct subset of longer conversations centred on professional development, likely related to CVs or performance reviews. These cases highlight the LLM’s use beyond core SOC functions, supporting analysts in broader professional contexts.
Conversations with four or more consecutive Command Understanding/Analysis steps suggest thorough exploration and careful interpretation of related commands in high-stakes environments. Other sequences chain distinct but related subtasks, combinations such as File Understanding $\rightarrow$ Command Understanding/Analysis $\rightarrow$ Threat Analysis $\rightarrow$ Summarise LLM Output indicate layered workflows: first clarifying data or command behaviour, then assessing threats, and finally seeking help with incident reporting. We also observed iterative development conversations with repeated cycles of Code Writing and Understanding, Understanding Errors, and Query Writing, suggesting that LLM supports implementation, error checking, refinement, and formatting. Several multi-step interactions focused on communication tasks—editing text, improving clarity, and crafting empathetic emails, combining structural edits with tone adjustments to produce polished, context-appropriate messages for diverse stakeholders.
Across these conversations, patterns suggest that SOC analysts often use LLMs as iterative, context-sensitive tools for reasoning, drafting, and refinement. Most conversations are short and focused, typically 2–3 steps, centred on tasks such as command understanding, text editing or coding tasks. While longer sessions are rare, they reveal more complex workflows, indicating that LLMs can support layered analytical and communicative goals. These findings suggest emerging usage patterns, though they should be interpreted with caution, given the predominance of brief exchanges and the absence of direct cognitive measures capturing analysts’ reasoning processes.
@@ 4.5 Analyst Level Analysis
Engagement with the LLM varied widely across analysts in query volume and thematic breadth. Analysts clustered into three groups: low-usage (Cluster 0, $n=34$), moderate (Cluster 1, $n=10$), and a single high-volume outlier (Cluster 2, $n=1$).
We examined daily LLM query activity across all analysts (Figure 6) and observed distinct engagement patterns. One analyst (A68f) consistently used the system at a high volume across multiple months. A group of moderate users (e.g., Ac9a, A5b3) engaged in concentrated bursts of activity, likely linked to specific investigations. In contrast, most analysts issued only a few queries on scattered days, indicating sporadic or task-specific use. These behavioural differences support cluster-based segmentation and highlight varied LLM integration styles.
@@
A30d and A5b3 mainly focused on Code, Script, and Regex Analysis, with moderate exploration of other tasks. Adf2 showed a balanced profile across Text Generation, Command Understanding, and Tool/System Understanding, using the LLM for both exploratory and functional support. A4df combined command and code tasks with moderate Query/Data Processing and File Analysis, indicating technically grounded use. Among lower-volume users, such as A180, Ac48, and A95e, task engagement was more narrowly focused, typically confined to one or two functional categories such as Text Editing or Command Understanding. Notably, the recurring appearance of Command Understanding and Text Generation across many analysts underscores their central role in LLM support within SOC workflows.
@@
We observed that LLMs are not used uniformly across analysts. Instead, analysts appear to adapt the tool to their task demands, specific needs, and investigative styles, ranging from quick clarification and script debugging to collaborative drafting and iterative refinement. The LLM successfully supported this diversity without requiring extensive prompt engineering. This highlights that a well-configured, domain-aligned LLM can support a broad range of task types and prompt styles, enabling quick and effective interaction, even without deep prompting expertise.
@@ 4.5.1 Discussion on Potential Disengagement Patterns
A2f7 repeatedly queried the model version (perhaps unmet transparency need), A5a6 received two consecutive erroneous answers, with a misclassification ending the session, and A6e9 disengaged immediately after the LLM refused a request. Another analyst (Ae46) ceased engagement after several command analysis queries, although no clear failure pattern was identified. Further analysis revealed that task theme rankings for this group mirrored the broader pool of analysts, indicating that early drop-off was not task-specific. Collectively, this group suggests that early disengagement occurred for multiple reasons: in some cases without clear frustration, and in others following specific model failures.
Group 2: Persistent but Mixed-Experience Analysts (n = 6) These analysts remained active for four to eight months. Two analysts (Ae32, Aae4) experienced apparent errors yet persisted, indicating that perceived utility can offset isolated failures. Other patterns suggest diverse usage motives: A9c2 moved past an early debugging issue; A214 relied on the LLM for repeated Sumo Logic refinements; A8bb may have engaged only when necessary; and A525 mixed non-SOC queries with SOC ones. This group suggests that analysts tolerate occasional mistakes, perhaps if using the LLM for specific tasks is beneficial. While these patterns are informative, future work should observe SOC analysts and conduct interviews to validate disengagement reasons.
@@ 5 Discussion
This study set out to investigate a core question (restated briefly): How do SOC analysts utilise LLMs in their daily workflows, including the specific tasks they apply them to, how these uses align with established cybersecurity frameworks (e.g., NICE), and the conversational patterns that characterise their interactions? Below, we address this question through a multi-level analysis and conclude with a discussion of the study’s limitations and the broader implications for designing LLM-based assistance in SOCs.
@@ 5.1 Query Level: Analyst Needs and Task Focus
The progressive rise in LLM use across tasks (Figure 6) and sustained engagement by more analysts (Figure 3) show that analysts gradually integrated the LLM into daily workflows, applying it to more tasks over time (see Figure 9 in Appendix). These usage patterns suggest that for some analysts, the LLM delivered sufficient value to merit repeated use during investigative sessions.
@@
Query patterns reflected clear cognitive intent. Analysts’ queries were not just syntactic variations but reflected clear cognitive intent. Functional queries like “what does [x] do” mapped to interpretive tasks such as command understanding, while requests like “reword this” reflected text refinement goals. This alignment between linguistic patterns and task types suggests that query structure can serve as a proxy for the analyst’s reasoning objective.
@@ 5.3 Analyst-Level: Usage Patterns and Integration Approaches
LLM use varies. Analysts differed not only in the frequency of LLM engagement but also in the types of tasks they delegated and how they framed their requests. Some used LLMs for command parsing and debugging; others focused on editing or summarising. These distinctions suggest an opportunity for investigating the effectiveness of role-aware, adaptive integration strategies that personalise tooling based on emerging usage patterns. Our findings build on early claims of LLMs as cognitive aides and validate emerging theories of diverse user strategies and interaction modes [7, 8, 13, 39]. While prior work explored LLMs for structured extraction and task acceleration, we show that usage patterns naturally align with analysts’ task orientations, without prompt engineering or adaptation.
LLMs function primarily as evidence interpreters, not decision-makers. Only 4% of queries sought explicit recommendations (e.g., “is this malicious?”). Instead, most analysts requested explanation, interpretation, or contextualisation, highlighting a strong preference for maintaining decision authority. Our findings support designing LLMs as evaluative aides that surface evidence without prescribing actions, aligning with concerns about over-reliance on automation and reinforcing the need for expert-centred, machine-in-the-loop systems [41, 43, 44, 45]. Analysts’ preference for interpretation over recommendations warrants further study.
5.4 Other factors that modulate routine engagement: isolated errors and trust.
@@
Research Opportunity: Future studies could investigate whether the preference for evidence-oriented outputs (4% in our study) generalises to different settings, and also compare the impact of evidence-oriented versus recommendation-oriented outputs on accuracy, response time, and calibrated trust [40, 41] to determine which support style minimises misuse yet maintains efficiency during high-pressure triage.